# Unlearning or Concealment? A Critical Analysis and Evaluation Metrics for Unlearning in Diffusion Models

### v1

@@ Abstract
Recent research has seen significant interest in methods for concept removal and targeted forgetting in diffusion models. In this paper, we conduct a comprehensive white-box analysis to expose significant vulnerabilities in existing diffusion model unlearning methods. We show that the objective functions used for unlearning in the existing methods lead to decoupling of the targeted concepts (meant to be forgotten) for the corresponding prompts. This is concealment and not actual unlearning, which was the original goal. Ideally, the knowledge of the concepts should have been completely scrubbed in the latent space of the model. The ineffectiveness of current methods stems primarily from their narrow focus on reducing generation probabilities for specific prompt sets, neglecting the diverse modalities of intermediate guidance employed during the inference process. The paper presents a rigorous theoretical and empirical examination of four commonly used techniques for unlearning in diffusion models, while exposing their potential weaknesses. We introduce two new evaluation metrics: Concept Retrieval Score ($\mathcal{CRS}$) and Concept Confidence Score ($\mathcal{CCS}$). These metrics are based on a successful adversarial attack setup that can recover forgotten concepts from unlearned diffusion models. $\mathcal{CRS}$ measures the similarity between the latent representations of the unlearned and fully trained models after unlearning. It reports the extent of retrieval of the forgotten concepts with increasing amount of guidance. $\mathcal{CCS}$ quantifies the confidence of the model in assigning the target concept to the manipulated data. It reports the probability of the unlearned model’s generations to be aligned with the original domain knowledge with increasing amount of guidance. The proposed metrics offer stringent assessment of unlearning in diffusion models, enabling a more accurate evaluation of concept erasure methods. Evaluating the existing unlearning methods with our metrics reveal significant shortcomings in their ability to truly unlearn concepts. Project page: https://respailab.github.io/unlearning-or-concealment
footnotetext: Corresponding author
@@ 1 Introduction
Diffusion models (Ho, Jain, and Abbeel 2020; Dhariwal and Nichol 2021; Kawar et al. 2022; Gu et al. 2023) have rapidly emerged as powerful tools for generating high-quality images and videos. However, their ability to generate content in an uncontrolled and unpredictable manner raises serious concerns regarding the misuse of these models. As a result, there has been growing interest in developing methods to unlearn or erase concepts from diffusion models (Kumari et al. 2023; Zhang et al. 2024; Heng and Soh 2024; Gandikota et al. 2023; Kim et al. 2023) to prevent the generation of harmful or undesired outputs.
@@
Recent approaches have explored various techniques for unlearning, targeting specific aspects of concept removal. For instance, (Gandikota et al. 2023) propose subtracting the prompt-conditioned noise prediction from the unconditional noise prediction, thereby leading the model away from generating the targeted concept. They introduce two variants: ESD-x, which fine-tunes the cross-attention layers for text-specific unlearning, and ESD-u, which fine-tunes unconditional layers for more generalized concept removal. Another approach by (Kumari et al. 2023) aims to overwrite the target concept by mapping it to an anchor concept distribution, though this method does not guarantee complete removal. Similarly, (Kim et al. 2023) perform self-distillation to align the conditional noise predictions of the targeted concept with their unconditional variants, enabling the erasure of multiple concepts simultaneously. Other works related to diffusion unlearning and machine unlearning in general include (Han et al. 2024; Yang et al. 2023; Fan et al. 2024; Tsai et al. 2023; Yang et al. 2024; Tarun et al. 2023b; Chundawat et al. 2023a, b; Tarun et al. 2023a; Hong, Lee, and Woo 2024).
Despite these advances, these unlearning methods have significant limitations. Most rely on regularization techniques or iterative refinement to remove targeted concepts from the model’s latent space. However, their objective functions often lead to decoupling targeted concepts from their corresponding prompts rather than achieving true concept erasure. This process effectively conceals the information rather than fully unlearning it, allowing hidden traces to resurface during generation. The primary flaw lies in focusing solely on reducing the generation probability for specific prompt sets while neglecting the varied modalities of intermediate guidance used during inference.
@@
Challenges in Evaluating Unlearning. Existing evaluation metrics for unlearning in diffusion models (Gandikota et al. 2023; Kumari et al. 2023; Kim et al. 2023; Lu et al. 2024; Gandikota et al. 2024; Heng and Soh 2024; Pham et al. 2023) generally focus on the final generated output, using metrics such as FID score, KID score, CLIP score (Hessel et al. 2021; Radford et al. 2021), and LPIPS. While these metrics assess the visual fidelity and prompt alignment of the output, they overlook the diffusion process’s latent stages. This leaves room for adversaries to introduce subtle modifications that can reinstate forgotten concepts during the generation pipeline. The discrepancy between perceived forgetting at the output level and the actual underlying model behavior highlights the inadequacy of current evaluation methods (Pham et al. 2023).
@@
Our Contributions. To address these challenges, we propose two new evaluation metrics designed to more accurately assess unlearning in diffusion models. Our approach focuses on the latent stages of the diffusion process, enabling a more comprehensive evaluation of concept erasure techniques. We provide a thorough theoretical and empirical analysis of these metrics, revealing the substantial limitations of existing methods when applied to four widely-used unlearning techniques in diffusion models. Our experimental results demonstrate the effectiveness of our proposed metrics, underscoring the need for a more critical and rigorous evaluation of unlearning methods in generative models.
@@
  • Comprehensive White-Box Analysis. We conduct an in-depth analysis of existing diffusion model unlearning methods, exposing significant vulnerabilities. Our study reveals that current techniques often lead to concept concealment rather than complete unlearning, leaving traces of targeted knowledge within the model’s latent space.
  • New Evaluation Metrics. We introduce two novel metrics—Concept Retrieval Score ($\mathcal{CRS}$) and Concept Confidence Score ($\mathcal{CCS}$)—that offer a more rigorous assessment of unlearning effectiveness. These metrics, rooted in an adversarial attack framework, measure the retrieval of supposedly forgotten concepts and the model’s confidence in generating related content.
  • Critical Assessment of Existing Methods. Our theoretical and empirical analysis of four common unlearning techniques reveals critical flaws in achieving true concept erasure. By applying our new metrics, we demonstrate the persistent limitations of these methods, advocating for more robust approaches to machine unlearning in generative models.
These contributions advance the understanding of machine unlearning in diffusion models and provide valuable tools for evaluating the effectiveness of unlearning techniques. Our work paves the way for more secure and controllable generative AI systems capable of genuine concept erasure.
@@ 2 Preliminaries
Diffusion Models. Denoising Diffusion Models (DDMs) generate images through a sequential denoising process that transforms an initial random Gaussian noise input into a coherent image. This iterative refinement operates over a series of discrete time steps. Latent Diffusion Models (LDMs) (Ho, Jain, and Abbeel 2020) enhance DDMs by performing this process within a reduced-dimensional latent space, leveraging an encoder-decoder architecture. The diffusion occurs in this latent space, directed by a neural network trained to model the denoising dynamics. This approach facilitates both unconditional and conditional image generation by modulating the latent representation according to specified conditions or prompts. The denoising process in LDMs is mathematically described by the following equation:
@@
General Equation for Diffusion Unlearning. The process of machine unlearning in diffusion models can be expressed through the following equation, which outlines how the model’s parameters are adjusted to remove specific information:
$\theta^{\text{unlearned}}=\theta^{\text{fully\_trained}}-\eta\cdot\sum_{t=1}^{T}\nabla_{\theta}\mathcal{L}_{t}(\theta;\mathcal{D}_{r},x_{t},c)$ (2)
where $\theta^{\text{unlearned}}$ is model parameter after the unlearning process, $\theta^{\text{fully\_trained}}$ is fully trained model parameter before unlearning. $\eta$ is the learning rate, $T$ is total number of diffusion steps, $c$ denotes the conditions imposed during unlearning and $\mathcal{L}_{t}(\theta;\mathcal{D}_{r},x_{t},c)$ is loss function at step $t$. Post-unlearning, the diffusion model should generate images free of excluded concepts while retaining high quality and diversity.
@@
where $P_{\theta^{\text{unlearned}}}(c_{f}\mid x_{t}+\delta_{t})$ is the probability of generating the concept $c_{f}$ at step $t$ given the adversarially perturbed latent state $x_{t}+\delta_{t}$. $\delta_{t}$ is an adversarial perturbation applied at step $t$ to test the robustness of unlearning.
@@
$P_{\theta^{\text{unlearned}}}(c_{r}\mid x_{t})\approx P_{\theta^{\text{fully\_trained}}}(c_{r}\mid x_{t})\quad\forall t\in[1,T]$ (4)
Where $P_{\theta^{\text{unlearned}}}(c_{r}\mid x_{t})$ is the probability of generating the concept $c_{r}$ at step $t$ after unlearning. $P_{\theta^{\text{fully\_trained}}}(c_{r}\mid x_{t})$ is the probability of generating the concept $c_{r}$ at step $t$ before unlearning.
@@
Standard metrics like FID, KID, CLIP score, and LPIPS assess image quality, diversity, and semantic alignment, but may not fully capture successful unlearning. This underscores the need for advanced metrics that specifically evaluate the removal of unlearned concepts, offering a deeper insight into the model’s performance after unlearning.
@@
We also provide a rigorous mathematical formulation of the unlearning process using optimal transport theory, specifically through Earth Mover’s Distance (EMD), to assess the effectiveness of unlearning in the Appendix
3 Probing the Effectiveness of Unlearning with Proposed Evaluation Metrics
Experimental setup. To comprehensively evaluate the effectiveness of the unlearning process and the model’s ability to retain or align with the undesired domain knowledge, we generate reference image sets that serve as benchmarks. These reference sets capture the original domain knowledge and the unlearned domain knowledge, enabling a direct comparison with the images generated during the partial diffusion process.
Partial Diffusion: We employ partial diffusion to selectively impart heavily noised features of the forgotten concept at a linear pace. This helps us ascertain whether the model can recall forgotten concepts after reintroducing a small fraction of its latent code. It involves dividing the denoising process into multiple stages or experts, each focusing on a specific slices of the denoising process. We deploy partial diffusion in two ways 1 The prompt is passed through the the fully trained model, which performs the initial stages of denoising, generating a partially denoised output based on a certain percentage of the total timestep $T$. 2 The partially denoised output from the fully trained model is then used as the input for the unlearned model, which takes over and completes the remaining denoising steps, producing the final output.
Using a prompt $\mathcal{P}$ that encompasses the unlearned concept and varying generation seeds, three distinct datasets are generated:
Unlearned Domain Knowledge ($\lambda_{\mathcal{U}}$): We generate this dataset using prompt $\mathbf{p}$ with the unlearned model ($\theta^{*}$) for $\lambda$ steps, representing the post-unlearning domain knowledge. These images serve as a reference for the desired unlearning outcome, reflecting the removed concept.
Original Domain Knowledge ($\lambda_{\mathcal{O}}$): This dataset is generated using prompt $\mathcal{P}$ with the original model ($\theta$) for $\lambda$ steps, representing pre-unlearning domain knowledge. These images serve as a reference for the concept to be unlearned.
Partially Diffused Knowledge ($\lambda_{\mathcal{P}}$): This set is generated using prompt $\mathbf{p}$ and a fixed seed, varying the partial diffusion ratio $\psi$. It comprises $\mathcal{N}$ images with unique $\psi$ values, representing potential leakage of unlearned knowledge from model $\theta^{*}$.
Our evaluation framework utilizes reference image sets to provide a comprehensive and unbiased assessment of the unlearning process. By comparing images generated with varying partial diffusion ratios to these reference sets, we quantify the model’s success in unlearning targeted concepts. The use of constant prompts and varying seeds ensures representative and fair evaluation across different models and parameters. The visual nature of these sets offers intuitive interpretability, allowing for qualitative assessment of the unlearning effectiveness. Combined with appropriate metrics, this approach forms a robust foundation for analyzing the model’s alignment with desired domain knowledge post-unlearning. The proposed partial diffusion pipeline is absolute in its nature and is not dependent on any particular form of modality. It simply provides a partially denoised latent along with an optional form of modality to the model, allowing it to ”continue” from where the first expert left off. The model predictions henceforth are absolute and will function in any diffusion-based model, regardless of its conditioning method.
@@ 3 Probing the Effectiveness of Unlearning with Proposed Evaluation Metrics
Figure 1: The proposed Partial Diffusion process to extract “forgotten” concepts from the unlearned model.
Concept Confidence Score($\mathcal{CCS}$.) The $\mathcal{CCS}$ utilizes a fine-tuned ResNet-18 model for binary classification to differentiate between original ($\lambda_{\mathcal{O}}$) and unlearned ($\lambda_{\mathcal{U}}$) domain knowledge. This model predicts the probability whether a generated image belongs to the original domain. Let $\lambda_{\mathcal{P}}=\{p_{1},p_{2},\ldots,p_{\mathcal{N}}\}$ be the set of images generated after partial diffusion, where each $p_{i}$ is an image. The probability that a generated image $p_{i}$ belongs to the original domain $\lambda_{\mathcal{O}}$ is denoted as $P(y=\lambda_{\mathcal{O}}\mid p_{i})$. The $\mathcal{CCS}$ for retaining the original domain knowledge is given as
$\mathcal{CCS}$ measures unlearning effectiveness in diffusion models by quantifying generated images’ association with original domain knowledge $\lambda_{\mathcal{O}}$. A low $\mathcal{CCS}_{\text{retain}}$ and a high $\mathcal{CCS}_{\text{forget}}$ indicate that the model has effectively erased the specified concepts while maintaining its generative capabilities.
Concept Retrieval Score($\mathcal{CRS}$.) The $\mathcal{CRS}$ is computed using cosine similarity between feature embeddings of generated images and ground truth images from original and unlearned domains. Let $\lambda_{\mathcal{P}}=\{p_{1},p_{2},\ldots,p_{\mathcal{N}}\}$ represent the set of partially diffused knowledge (i.e., generated images), $\lambda_{\mathcal{O}}=\{o_{1},o_{2},\ldots,o_{\lambda}\}$ be the set of images from the original domain knowledge, and $\lambda_{\mathcal{U}}=\{u_{1},u_{2},\ldots,u_{\lambda}\}$ be the set of images from the unlearned domain knowledge. The feature embeddings for these images are extracted using a fine-tuned ResNet-18 model. We denote the feature embeddings for generated images, original domain images, and unlearned domain images by $f(p_{i})$, $f(o_{i})$, and $f(u_{i})$, respectively. The $\mathcal{CRS}$ for retaining the original domain knowledge is computed as
The $\mathcal{CRS}$ for unlearning the targeted concept is calculated as
The terms $\cos(f(p_{i}),f(o_{i}))$ and $\cos(f(p_{i}),f(u_{i}))$ represent the cosine similarities between the feature embeddings of the generated image $p_{i}$ with the original domain image $o_{i}$ and the unlearned domain image $u_{i}$, respectively. The $arctangent$ function scales these similarities to a meaningful range for better interpretation. The $\mathcal{CRS}$ quantifies the alignment between generated images and the original or unlearned domain knowledge, measured through cosine similarity of feature embeddings extracted from a fine-tuned ResNet-18 model. A high $\mathcal{CRS}_{\text{forget}}$ indicates effective unlearning, as it shows reduced similarity to the original domain, while a low $\mathcal{CRS}_{\text{retain}}$ suggests generated images remain closely aligned with the original domain knowledge.

### v2

@@ Abstract
Recent research has seen significant interest in methods for concept removal and targeted forgetting in text-to-image diffusion models. In this paper, we conduct a comprehensive white-box analysis showing the vulnerabilities in existing diffusion model unlearning methods. We show that existing unlearning methods lead to decoupling of the targeted concepts (meant to be forgotten) for the corresponding prompts. This is concealment and not actual forgetting, which was the original goal. The targeted concepts remain embedded in the model’s latent space, allowing them to be generated. Current methods are ineffective mainly because they focus too narrowly on lowering generation probabilities for certain prompts, overlooking the different types of guidance used during inference. This paper presents a rigorous theoretical and empirical examination of four commonly used techniques for unlearning in diffusion models, while showing their potential weaknesses. We introduce two new evaluation metrics: Concept Retrieval Score ($\mathcal{CRS}$) and Concept Confidence Score ($\mathcal{CCS}$). These metrics are based on a successful adversarial attack setup that can recover forgotten concepts from unlearned diffusion models. $\mathcal{CRS}$ measures the similarity between the latent representations of the unlearned and fully trained models after unlearning. It reports the extent of retrieval of the forgotten concepts with increasing amount of guidance. $\mathcal{CCS}$ quantifies the confidence of the model in assigning the target concept to the manipulated data. It reports the probability of the unlearned model’s generations to be aligned with the original domain knowledge with increasing amount of guidance. The $\mathcal{CCS}$ and $\mathcal{CRS}$ enable a more robust evaluation of concept erasure methods. Evaluating existing five state-of-the-art methods with our metrics, reveal significant shortcomings in their ability to truly unlearn. Source Code: https://respailab.github.io/unlearning-or-concealment
**footnotetext: Corresponding author
@@ 1 Introduction
Diffusion models [20, 7, 23, 14] have rapidly emerged as powerful tools for generating high-quality images and videos. However, their ability to generate content in an uncontrolled and unpredictable manner raises serious concerns regarding the misuse of these models. As a result, there has been growing interest in developing methods to regulate with unlearning or erasing concepts from diffusion models [25, 52, 17, 12, 24] to prevent the generation of harmful or undesired outputs.
@@
Recent unlearning approaches target specific aspects of concept removal. For example, [12] subtracts prompt-conditioned noise from unconditional noise predictions, guiding the model away from generating the targeted concept. The two variants include ESD-x: fine-tuning cross-attention layers for text-specific unlearning, and ESD-u: fine-tunes unconditional layers for broader concept removal. Another method [25] attempts to overwrite the target concept by mapping it to an anchor distribution, though it doesn’t ensure complete removal. [24] perform self-distillation to align the conditional noise predictions of the targeted concept with their unconditional variants, enabling the erasure of multiple concepts simultaneously. Other works related to diffusion unlearning and machine unlearning in general include [16, 49, 9, 45, 50, 44, 5, 6, 43, 21, 27]. These unlearning methods rely on regularization techniques or iterative refinement to remove targeted concepts from the model’s latent space. However, their objective functions tend to decouple targeted concepts from associated prompts rather than achieving genuine concept erasure. This approach often obscures, rather than fully unlearns, the information, allowing hidden traces to re-emerge during generation. A key issue is the narrow focus on reducing generation probability for specific prompt sets, which overlooks the diverse types of intermediate guidance employed throughout the inference process.
@@
Limitations in existing evaluation metrics for unlearning in text-to-image diffusion models. Existing evaluation metrics for unlearning in diffusion models [12, 25, 24, 29, 13, 17, 32] generally focus on the final generated output, using metrics such as FID score, KID score, CLIP score [18, 34], and LPIPS. While these metrics assess the visual fidelity and prompt alignment of the output, they overlook the diffusion process’s latent stages. This leaves room for adversaries to introduce subtle modifications that can reinstate forgotten concepts during the generation pipeline. The discrepancy between perceived forgetting at the output level and the actual underlying model behavior highlights the inadequacy of current evaluation methods [32].
@@
Our contributions. To address these challenges, we propose two new evaluation metrics designed to more robustly assess unlearning in diffusion models. Our approach focuses on the latent stages of the diffusion process, enabling a more comprehensive evaluation of concept erasure techniques. We provide a thorough theoretical and empirical analysis of these metrics, revealing the substantial limitations of existing methods when applied to five widely-used unlearning techniques. Our experimental results demonstrate the effectiveness of our proposed metrics, underscoring the need for a more critical and rigorous evaluation of unlearning methods in generative models.
@@
  • New evaluation metrics. We introduce two new metrics—Concept Retrieval Score ($\mathcal{CRS}$) and Concept Confidence Score ($\mathcal{CCS}$)—that offer a more rigorous assessment of unlearning effectiveness. These metrics, rooted in an adversarial attack framework, measure the retrieval of supposedly forgotten concepts and the model’s confidence in generating related content.
  • White-box analysis of existing methods. We conduct an in-depth analysis of existing unlearning methods for diffusion models, revealing their vulnerabilities. Our findings show that current techniques often result in concept concealment rather than complete unlearning, leaving residual traces of targeted knowledge that can still generate the forgotten concepts.
  • Comparative analysis with existing metrics. We present a comparative analysis of our metrics alongside established metrics like KID and CLIP scores. This analysis highlights the need for more robust evaluation methods for machine unlearning in generative models.
@@ 2 Preliminaries
Diffusion models. Denoising Diffusion Models (DDMs) generate images through a sequential denoising process that transforms an initial random Gaussian noise input into a coherent image. This iterative refinement operates over a series of discrete time steps. Latent Diffusion Models (LDMs) [20] enhance DDMs by performing this process within a reduced-dimensional latent space, leveraging an encoder-decoder architecture. The diffusion occurs in this latent space, directed by a neural network trained to model the denoising dynamics. This approach facilitates both unconditional and conditional image generation by modulating the latent representation according to specified conditions or prompts. The denoising process in LDMs is mathematically described by the following equation:
@@
$P_{\theta^{\text{u}}}(c_{f}\mid x_{t}+\delta_{t})\approx 0\quad\forall t\in[1,T]$ (2)
where $P_{\theta^{\text{u}}}(c_{f}\mid x_{t}+\delta_{t})$ is the probability of generating the concept $c_{f}$ at step $t$ given the adversarially perturbed latent state $x_{t}+\delta_{t}$. $\theta^{\text{u}}$ is model parameter after the unlearning, $\delta_{t}$ is an adversarial perturbation applied at step $t$ to test the robustness of unlearning.
@@
$P_{\theta^{\text{u}}}(c_{r}\mid x_{t})\approx P_{\theta^{\text{o}}}(c_{r}\mid x_{t})\quad\forall t\in[1,T]$ (3)
@@
where $P_{\theta^{\text{u}}}(c_{r}\mid x_{t})$ is the probability of generating the concept $c_{r}$ at step $t$ after unlearning. $\theta^{\text{o}}$ is originally trained model. $P_{\theta^{\text{o}}}(c_{r}\mid x_{t})$ is the probability of generating the concept $c_{r}$ at step $t$ before unlearning. Existing standard metrics like FID, KID, CLIP score, and LPIPS assess visual fidelity and prompt alignment but overlook latent stages of the diffusion process, allowing adversaries to subtly reinstate forgotten concepts during generation. This underscores the need for advanced metrics that specifically evaluate the removal of unlearned concepts, offering a deeper insight into the model’s performance after unlearning.
@@
We also provide a rigorous mathematical formulation of the unlearning process using optimal transport theory, specifically through Earth Mover’s Distance (EMD), to assess the effectiveness of unlearning in LABEL:sec:emd_theory.
@@
Figure 1: The proposed partial diffusion process to extract forgotten concepts from the unlearned model.
3 Proposed Evaluation Metrics
3.1 Evaluation framework
To comprehensively evaluate the effectiveness of the unlearning process and the model’s ability to retain or align with the undesired domain knowledge, we generate reference image sets that serve as benchmarks. These reference sets capture the original domain knowledge and the unlearned domain knowledge, enabling a direct comparison with the images generated during the partial diffusion process.
Partial diffusion. We employ partial diffusion to selectively impart heavily noised features of the forgotten concept at a linear pace. This helps us ascertain whether the model can recall forgotten concepts after reintroducing a small fraction of its latent code. It involves dividing the denoising process into multiple stages or experts, each focusing on a specific slices of the denoising process. We deploy partial diffusion in two ways 1 The prompt is passed through the the fully trained model, which performs the initial stages of denoising, generating a partially denoised output based on a certain percentage of the total timestep $T$. 2 The partially denoised output from the fully trained model is then used as the input for the unlearned model, which takes over and completes the remaining denoising steps, producing the final output. The process if visual depicted in Figure 1. Using a prompt $\mathcal{P}$ that encompasses the unlearned concept and varying generation seeds, three distinct datasets are generated:
Unlearned domain knowledge ($\lambda_{\mathcal{U}}$): We generate this dataset using prompt $\mathbf{p}$ with the unlearned model ($\theta^{u}$) for $\lambda$ steps, representing the post-unlearning domain knowledge. These images serve as a reference for the desired unlearning outcome, reflecting the removed concept.
Original domain knowledge ($\lambda_{\mathcal{O}}$): This dataset is generated using prompt $\mathcal{P}$ with the original model ($\theta^{o}$) for $\lambda$ steps, representing pre-unlearning domain knowledge. These images serve as a reference for the concept to be unlearned.
Partially diffused knowledge ($\lambda_{\mathcal{P}}$): This set is generated using prompt $\mathbf{p}$ and a fixed seed, varying the partial diffusion ratio $\psi$. It comprises $\mathcal{N}$ images with unique $\psi$ values, representing potential leakage of unlearned knowledge from model $\theta^{u}$.
The step-by-step process of the partial diffusion pipeline is outlined in Algorithm 1.
Usability across different text-to-image models. Our evaluation framework utilizes reference image sets to provide an unbiased assessment of the unlearning process. By comparing images generated with varying partial diffusion ratios with these reference sets, we quantify the model’s success in unlearning targeted concepts. The use of constant prompts and varying seeds ensures representative and fair evaluation across different models and parameters. The visual nature of these sets offers intuitive interpretability, allowing for qualitative assessment of the unlearning effectiveness. Combined with appropriate metrics, this approach forms a robust foundation for analyzing the model’s alignment with desired domain knowledge post-unlearning. The proposed partial diffusion pipeline operates independently of any specific modality, offering a partially denoised latent with an optional modality input to guide the model from where the initial expert left off. Subsequent model predictions remain stable and apply universally across diffusion-based models, regardless of conditioning methods, ensuring flexibility and consistency in unlearning tasks.